# A Spectral Localization Method for Time-Fractional Integro-Differential Equations with Nonsmooth Data

### v1

@@ 2.3 Spatial Semi-discrete Galerkin Scheme
We consider a family of quasi-uniform triangulations ${\left\{\mathcal{T}_{h}\right\}}_{0<h<1}$ of the domain $\overline{\Omega}$, with $\bar{h}$ denoting the maximum diameter among all finite elements. On $\mathcal{T}_{h}$, construct the finite element function space $S_{h}:=S_{h}(\Omega)\subset H_{0}^{1}(\Omega)$ consisting of continuous, piecewise linear functions:
$S_{h}:=\left\{\chi\in H_{0}^{1}(\Omega):\chi|_{\tau}\text{ is a linear polynomial over }\tau\in\mathcal{T}_{j}\right\}.$
The semi-discrete Galerkin finite element approximation to problem (1.1) is then formulated as follows: find $u_{h}(t)\in S_{h}$ such that for $0<t\leq T,$
$\langle^{C}_{0}D_{t}^{\beta}u_{h},v_{h}\rangle+((\nabla u_{h}(t),\nabla v_{h})+\int_{0}^{t}\kappa_{\alpha}(t-s)(\nabla u_{h}(s),\nabla v_{h})\,ds$ $=\langle f,v_{h}\rangle\quad\forall v_{h}\in S_{h},$ (2.11)
$u_{h}(0)$ $=u_{0,h}.$
Here $u_{0,h}:=P_{h}u_{0}$ and $f_{h}(t):=P_{h}f(t)$, which denote the $L^{2}$-projections of $u_{0,h}$ and $f_{h}(t)$ on $S_{h}$, respectively. That is,
$P_{h}:L^{2}(\Omega)\to S_{h},\,(P_{h}u,v_{h})=(u,v_{h})\quad\forall v_{h}\in S_{h},u\in L^{2}(\Omega).$
For convenience, we rewirte the semidiscrete scheme (2.11) in weak form as
$^{C}_{0}D_{t}^{\beta}u_{h}+A_{h}u_{h}+\int_{0}^{t}\kappa_{\alpha}(t-s)A_{h}u_{h}(s)\,ds=f_{h}\quad\forall t>0,$ (2.12)
where $A_{h}$ is the Ritz projection operator [2] of $A$. After applying the Laplace transform to both sides of Eq. (2.12), we have
@@
$u_{h}(t)=\frac{1}{2\pi i}\int_{\Gamma}e^{zt}(m(z)I+A_{h})^{-1}\left(\frac{m(z)}{z}u_{0,h}+\frac{m(z)}{z^{\beta}}\widehat{f}_{h}(z)\right)\,dz,$ (2.14)
@@ 2.4 Fully Discrete Scheme and Acceleration of Algorithm
Applying the CIM to the semi-discrete scheme (2.14), the fully discrete scheme for problem (1.1) can be obtained as
$u_{h}^{N}(t):=I_{\tau;h;N}=\frac{\tau}{\pi}\text{Im}\left\{\sum_{k=0}^{N-1}e^{z_{k}t}\widehat{u}_{h}(z_{k})z_{k}^{\prime}\right\},$ (2.15)
where $\widehat{u}_{h}(z_{k})$ is obtained by solving Eq. (2.13) at $z=z_{k},k=0,1,2,...,N-1$.
From Eq. (2.15), it is evident that the solution at a given time $t=t_{n}$ can be computed without relying on any information from previous time. This approach, referred to as a localization method, provides considerable computational advantages over nonlocal schemes.
However, achieving optimal error convergence with the scheme (2.15) typically requires $N$ to be between 100 and 200, which envolves solving 100 to 200 elliptic equations (2.13). The computational cost remains substantial as the spatial dimension increases. To further reduce the computational cost, one may emply the barycentric Lagrange interpolation [1] to approximate all values of $\widehat{u}_{h}(z_{k})$, leveraging the independence among these computations. Specifically, we first select $n+1$ Chebyshev interpolation nodes $z_{x_{j}}$ for $j=0,1,\ldots,n,$ along the truncated horizontal line. Then the values of $\widehat{u}_{h}(z_{x_{j}})$ can be computed by solving the associated elliptic equations (2.13). Finally, the values of $\widehat{u}_{h}(z_{k})$ at the original equispaced nodes are approximated by using of the barycentric interpolation formula:
@@
In this manner, only $n+1$ elliptic equations need to be solved. Empirical results suggest that $n$ typically lies between $10$ and $20$. In view of the focus of this paper on localization and non-smoothness, acceleration algorithms will not be further discussed in subsequent sections. For further details on the implementation and error analysis of this accelerated algorithm, we refer the reader to Section $6$ of [20].
@@ 3.1 Solution Theory
and assume that $\widehat{f}(z)$ is analytic and bounded on $S\subseteq\Sigma_{\tilde{\vartheta}}$ for some $\tilde{\vartheta}\in(\pi/2,\pi)$, the associated norm is defined as
$\|\widehat{f}(z)\|_{S}:=\sup_{z\in S}\|\widehat{f}(z)\|_{L^{2}(\Omega)},\quad S\subseteq\Sigma_{\tilde{\vartheta}}$.
@@
In [22], Mahata and Sinha provide some important regularity properties for the solution $u(t)$ as follows:
@@
Let $j\in\{0,1\}$, $m\in\mathbb{N}$, and denote $u^{(m)}(t):=\frac{\partial^{m}u(t)}{\partial t^{m}},f^{(m)}(t):=\frac{\partial^{m}f(t)}{\partial t^{m}}$. For the solution $u(t)$ of problem (1.1), there are
@@
    (Nonhomogeneous Problem) If $u_{0}=0$ and $f\in C^{m}([0,T];L^{2}(\Omega))$, the weak solution $u(t)$ is unique, the following a priori estimates hold for positive time:
@@
The bound of $m(z)$ (see Eq. (2.7)) is an important tool in error analysis. In Lemma 2.2 of [22], Mahata and Sinha established an estimate for $|m(z)|$ under the condition $\alpha+\beta\geq 1$. This paper extends their result to a broader range $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ with any $0<\varepsilon<1$, leading to the following lemma:
@@
Let $m(z)$ be defined as in Eq. (2.7) with $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ for any fixed $0<\varepsilon<1$. Define $\tilde{\vartheta}:=\frac{(\alpha+\beta+\varepsilon)\pi}{2(\alpha+\beta)}$, and let $z\in\Sigma_{\tilde{\vartheta}}$. Then for any $z\in\Sigma_{\tilde{\vartheta}}$, there hold
@@
    $m(z)\in\Sigma_{\zeta}$ with $\zeta=\max{\left\{\left(\frac{\beta}{\alpha+\beta}\right)\cdot\frac{(\alpha+\beta+\varepsilon)\pi}{2},\frac{(\alpha+\beta+\varepsilon)\pi}{2}\right\}}$,
@@
Consequently, the resolvent satisfies the estimate:
$\|(m(z)I+A)^{-1}\|_{L^{2}(\Omega)}\leq C|m(z)|^{-1},\quad z\in\Sigma_{\tilde{\vartheta}}$. (3.1)
@@
$|\arg(m_{1})|=\beta|\psi|<\left(\frac{\beta}{\alpha+\beta}\right)\cdot\frac{(\alpha+\beta+\varepsilon)\pi}{2},$
@@
Hence, $m(z)\in\Sigma_{\zeta}$ with $\zeta=\max{\left\{\left(\frac{\beta}{\alpha+\beta}\right)\cdot\frac{(\alpha+\beta+\varepsilon)\pi}{2},\frac{(\alpha+\beta+\varepsilon)\pi}{2}\right\}}$. ∎
@@
Since $\beta$ is strictly positive, the conditions in Lemma 2 imply that $\Im(m(z))=0$ cannot occur. As a result, $(m(z)I+A)^{-1}$ is nonsingular, whereas $\widehat{u}(z)$ has a single singularity at $z=0$.
@@
Now, the determination of parameter $\tilde{c}$, a question left open in Section 2, can also be addressed. As indicated by Eq. (2.9), the slope of the asymptote of the hyperbolic contour $\Gamma$ is given by $\frac{\sin(\theta+c)}{\cos(\theta+c)}=\cot(\theta+c),$ that is, $\tan\delta=\cot(\theta+c),$ where $\delta$, as introduced earlier, denotes the angle between this asymptote and the real axis. It follows that
@@
Therefore, we have
$\tilde{c}=\min\left\{\theta,\frac{\pi}{2}-\theta,\frac{\varepsilon\pi}{2(\alpha+\beta)}-\theta\right\}.$
In fact, since $\alpha+\beta\in(\varepsilon,2-\varepsilon)$, then $\frac{\varepsilon\pi}{2(\alpha+\beta)}-\theta<\frac{\pi}{2}-\theta$. Thus, actually
$\tilde{c}=\min\left\{\theta,\frac{\varepsilon\pi}{2(\alpha+\beta)}-\theta\right\}.$
@@ 3.2 Error Analysis in Temporal Semi-discrete CIM Scheme
The error of the CIM scheme (2.10) can be decomposed into
$\mathrm{Err}_{\tau}:=\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}\leq\|I-I_{\tau}\|_{L^{2}(\Omega)}+\|I_{\tau}-I_{\tau;N}\|_{L^{2}(\Omega)},$
here the definitions of $I$, $I_{\tau}$ and $I_{\tau;N}$ are given by Eq. (2.3) and Eq. (2.5), respectively. Let $DE:=\left\|I-I_{\tau}\right\|_{L^{2}(\Omega)}$ denote the discretization error, and $TE:=\left\|I_{\tau}-I_{\tau;N}\right\|_{L^{2}(\Omega)}$ the truncation error. In [30], the authors analyze these two types of errors by considering the function the function $g(\phi,t)$ in Eq. (2.3) as complex-valued, i.e., $g(\nu+ic,t)$, where $g(\phi,t)$ admits an analytic continuation w.r.t. $\phi=\nu+ic$, and the time $t$ can be regarded as a given parameter. They assume that
[…]
which leads to the following estimate for the discretization error:
$DE\leq\frac{M(\tilde{c},t)}{e^{2\pi\tilde{c}/\tau}-1},$ (3.2)
implying that $DE=O(e^{-2\pi\tilde{c}/\tau})$. Estimate (3.2) indicates that larger values of $\tilde{c}$ result in smaller discretization errors and faster convergence.
As for the truncation error $TE$, assuming that $g(\phi,t)$ decays rapidly as $\phi=\nu+ic$, with $\nu\to\pm\infty$, it can be approximated by the magnitude of the last retained term, That is, for fixed $\tau$ and as $N\to\infty$,
$TE=O(\left|g(\tau N,t)\right|).$
Now we analyze the decay properties of $g(\phi,t)$.
@@
Let $g(\phi,t)$ be defined as in Eq. (2.4), where $z(\phi)\in S\subseteq\Sigma_{\tilde{\vartheta}}$, $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ with $0<\varepsilon<1$, and $\phi=\nu+ic\in K$, $|c|<\tilde{c}$. For $t>0$, if $u_{0}\in L^{2}(\Omega)$, and $\|\widehat{f}(z)\|_{S}<\infty$, then there exists a constant $C>0$ such that
$\|g(\phi,t)\|_{L^{2}(\Omega)}\leq C\varphi(\theta,\tilde{c})e^{\mu t}\left(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\right)e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu},$ (3.3)
@@
Firstly, we estimate $g(\phi,t)$ on the upper part of the domain $K$: $\phi_{+}=\nu+ic\in K$, $0<c<\tilde{c}$. From Eq. (2.4) and (2.8), we obtain
$g(\nu+ic,t)=\frac{\mu}{2\pi}e^{\mu t\{1-\sin(\theta+c-i\nu)\}}\cos(\theta+c-i\nu)\widehat{u}\bigl{(}\mu\{1-\sin(\theta+c-i\nu)\}\bigr{)}.$
Let $l^{\prime}=\theta+c<\theta+\tilde{c}$. Then the following estimate holds:
$\|g(\nu+ic,t)\|_{L^{2}(\Omega)}\leq Ce^{\mu t(1-\sin l^{\prime}\cosh\nu)}|\cos(l^{\prime}-i\nu)|\cdot\bigl{\|}\widehat{u}\bigl{(}\mu\{1-\sin(l^{\prime}-i\nu)\}\bigr{)}\bigr{\|}_{L^{2}(\Omega)}.$ (3.4)
The main task now is to estimate $\widehat{u}(z)$. Using the expression of $\widehat{u}(z)$ in Eq. (2.6) and the a prior estimate (3.1), it follows that for all $z\in\Sigma_{\tilde{\vartheta}}$ with $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ and $0<\varepsilon<1$,
$\|\widehat{u}(z)\|_{L^{2}(\Omega)}$ $=\bigl{\|}(m(z)I+A)^{-1}\bigl{(}\frac{m(z)}{z}u_{0}+\frac{m(z)}{z^{\beta}}\widehat{f}(z)\bigr{)}\bigr{\|}_{L^{2}(\Omega)}$
@@
$|\cos(l^{\prime}-i\nu)|\cdot\bigl{\|}\widehat{u}\bigl{(}\mu\{1-\sin(l^{\prime}-i\nu)\}\bigr{)}\bigr{\|}_{L^{2}(\Omega)}$ (3.5)
$\leq$ $C\frac{|\cos(l^{\prime}-i\nu)|}{|\mu(1-\sin(l^{\prime}-i\nu))|}\|u_{0}\|_{L^{2}(\Omega)}+C\frac{|\cos(l^{\prime}-i\nu)|}{|\mu(1-\sin(l^{\prime}-i\nu))|^{\beta}}\|\widehat{f}(z)\|_{L^{2}(\Omega)}$
$\leq$ $C\sqrt{\frac{1+\sin l^{\prime}}{1-\sin l^{\prime}}}\|u_{0}\|_{L^{2}(\Omega)}+C(|1-\sin(l^{\prime}-i\nu)|^{1-\beta})\sqrt{\frac{1+\sin l^{\prime}}{1-\sin l^{\prime}}}\|\widehat{f}(z)\|_{L^{2}(\Omega)},$
where in the second inequality of (3.5), we have used the result from [17]:
$\left|\frac{\cos(l^{\prime}-i\nu)}{(1-\sin(l^{\prime}-i\nu))}\right|^{2}=\frac{\cosh\nu+\sin l^{\prime}}{\cosh\nu-\sin l^{\prime}}\leq\frac{1+\sin l^{\prime}}{1-\sin l^{\prime}}.$
Given that $0<\beta<1$, it follows that $|1-\sin(l^{\prime}-i\nu)|^{1-\beta}\leq 2^{1-\beta}$. Then according to Eq. (3.4) and (3.5), there holds
$\|g(\nu+ic,t)\|_{L^{2}(\Omega)}\leq C\sqrt{\frac{1+\sin(\theta+\tilde{c})}{1-\sin(\theta+\tilde{c})}}e^{\mu t}\left(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\right)e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu}.$
@@
$\|g(\nu-ic,t)\|_{L^{2}(\Omega)}\leq C\sqrt{\frac{1+\sin(\theta+\tilde{c})}{1-\sin(\theta+\tilde{c})}}e^{\mu t}\left(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\right)e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu}.$
@@
From Theorem 1, it can be observed that $g(\nu+ic,t)$ exhibits biexponential decay w.r.t $\nu$, that is, it decays quite rapidly as $\phi=\nu+ic$, $\nu\to\pm\infty$. To further analyze the error of the CIM, we need to introduce the following lemma.
@@
$\int_{l}^{+\infty}e^{-\gamma\cosh x}\,dx\leq\bigl{(}1+L(\gamma)\bigr{)}e^{-\gamma\cosh l},\quad l>0,\gamma>0.$
The error analysis of the CIM is presented in the following theorem. As noted in [37], each new value of $t$ leads to different parameters $\theta$ and $\tilde{c}$, implying that the transform $\widehat{u}(z)$ must be sampled on a distinct contour. This could result in computational inefficiency in the matrix problem associated with the discretization of the spatial operator $A$ in Eq. (2.13), as well as in the observation of the long-term dynamic evolution of the simulated solution over time. To address this and select a unified optimal contour $\Gamma$, a scaling parameter $\Lambda>1$ is introduced. The subsequent analysis is performed over the interval $t\in[t_{0},t_{1}]$, where $t_{1}=\Lambda t_{0}$.
@@
Let $u(t)$ and $u^{N}(t)$ be the solutions given by (2.2) and (2.10), respectively. Consider the time interval $t\in[t_{0},t_{1}]$ with $t_{1}=\Lambda t_{0}$ for some $\Lambda\geq 1$. Let $\Gamma$ be the contour defined in (2.8) with parameters $\theta>0,\mu>0$, and let $z\in\Sigma_{\tilde{\vartheta}}$ with $\alpha+\beta\in(\varepsilon,2-\varepsilon)$, $0<\varepsilon<1$. Then, the following error estimate holds uniformly:
$\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}\leq C\varphi(\theta,\tilde{c})L\bigr{(}\mu t_{0}\sin(\theta-\tilde{c})\bigl{)}\left(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\right)\times\left(\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\right),$

### v2

@@ 2.3 Spatial Semi-discrete Galerkin Scheme
We consider a family of quasi-uniform triangulations ${\{\mathcal{T}_{h}\}}_{0<h<1}$ of the domain $\bar{\Omega}$, with $\bar{h}$ denoting the maximum diameter among all finite elements, and construct the piecewise linear finite element function space $X_{h}\subset H_{0}^{1}(\Omega)$ over $\mathcal{T}_{h}$ consisting of continuous, piecewise linear functions:
$X_{h}:=\big\{\chi\in H_{0}^{1}(\Omega):\chi|_{\tau}\text{ is a linear polynomial over }\tau\in\mathcal{T}_{h}\big\}.$
The semi-discrete Galerkin finite element approximation to problem (1.1) is then formulated as follows: find $u_{h}(t)\in X_{h}$ such that for $0<t\leq T,$
$\left\{\begin{aligned} &\big({{}_{0}^{C}}\!D_{t}^{\beta}u_{h}(t),v_{h}\big)+\big(\nabla u_{h}(t),\nabla v_{h}\big)+\int_{0}^{t}\kappa_{\alpha}(t-s)\big(\nabla u_{h}(s),\nabla v_{h}\big)\,ds=(f,v_{h})\quad\forall v_{h}\in X_{h},\\ &u_{h}(0)=u_{0,h}.\end{aligned}\right$. (2.13)
Here $u_{0,h}:=P_{h}u_{0}$ denotes the $L^{2}$-projections of $u_{0}$ on $X_{h}$, where the $L^{2}$-projection operator $P_{h}$ is defined by
$P_{h}:L^{2}(\Omega)\to X_{h},\,(P_{h}u,v_{h})=(u,v_{h})\quad\forall v_{h}\in X_{h},u\in L^{2}(\Omega).$
For convenience, we rewrite the semidiscrete scheme (2.13) in weak form as
${{}_{0}^{C}}\!D_{t}^{\beta}u_{h}+A_{h}u_{h}+\int_{0}^{t}\kappa_{\alpha}(t-s)A_{h}u_{h}(s)\,ds=f_{h}\quad\forall t>0,$ (2.14)
where $f_{h}(t):=P_{h}f(t)$ and $A_{h}$ is the Ritz projection operator [3] of $A$. After applying the Laplace transform to both sides of (2.14), we have
@@
$u_{h}(t)=\frac{1}{2\pi i}\int_{\Gamma}e^{zt}(m(z)I+A_{h})^{-1}\Big(\frac{m(z)}{z}u_{0,h}+\frac{m(z)}{z^{\beta}}\widehat{f}_{h}(z)\Big)dz,$ (2.16)
@@ 2.4 Fully Discrete Scheme and Acceleration of Algorithm
Applying the CIM to the spatial semi-discrete scheme (2.16), we can obtain the fully discrete scheme for problem (1.1) as
$u_{h}^{N}(t):=I_{\tau;h;N}=\frac{\tau}{\pi}\text{Im}\Big\{\sum_{k=0}^{N-1}e^{z_{k}t}\widehat{u}_{h}(z_{k})z_{k}^{\prime}\Big\},$ (2.17)
where $\widehat{u}_{h}(z_{k})$ is obtained by solving (2.15) at $z=z_{k},k=0,1,2,...,N-1$.
From (2.17), it is evident that the solution at a given time $t$ can be computed without relying on any information from previous time. This approach, referred to as a localization method, provides considerable computational advantages over nonlocal schemes.
However, achieving optimal error convergence with the scheme (2.17) typically requires $N$ to be between 100 and 200, which involves solving 100 to 200 elliptic equations (2.15). The computational cost remains substantial as the spatial dimension increases. To further reduce the computational cost, one may employ the barycentric Lagrange interpolation [2] to approximate all values of $\widehat{u}_{h}(z_{k})$, leveraging the independence among these computations. Specifically, we first select $n+1$ Chebyshev interpolation nodes $z_{x_{j}}$ for $j=0,1,\ldots,n,$ along the truncated horizontal line. Then the values of $\widehat{u}_{h}(z_{x_{j}})$ can be computed by solving the associated elliptic equations (2.15). Finally, the values of $\widehat{u}_{h}(z_{k})$ at the original equispaced nodes are approximated by using of the barycentric interpolation formula:
@@
In this manner, only $n+1$ elliptic equations need to be solved. Empirical results suggest that $n$ typically lies between $10$ and $20$, thereby enabling a further notable reduction in computational load. In view of the focus of this paper on localization and nonsmoothness, acceleration algorithms will not be further discussed in subsequent sections. For further details on the implementation and error analysis of this accelerated algorithm, we refer the reader to [20, Section 6].
@@ 3.1 Solution Theory
and assume that $\widehat{f}(z)$ is analytic and bounded on $S\subseteq\Sigma_{\tilde{\vartheta}}$ (the domain $S$ is shown in Figure 2) for some $\tilde{\vartheta}\in(\pi/2,\pi)$, the associated norm is defined as
$\|\widehat{f}(z)\|_{S}:=\sup_{z\in S}\|\widehat{f}(z)\|_{L^{2}(\Omega)},\quad S\subseteq\Sigma_{\tilde{\vartheta}}$. (3.1)
@@
In [22], Mahata and Sinha provided some important regularity properties for the solution $u(t)$ as follows:
@@
Let $j\in\{0,1\}$, $m\in\mathbb{N}$, and denote $u^{(m)}(t):=\frac{\partial^{m}u(t)}{\partial t^{m}},f^{(m)}(t):=\frac{\partial^{m}f(t)}{\partial t^{m}}$. The solution $u(t)$ of problem (1.1) satisfies the following properties.
@@
    (Nonhomogeneous Problem) If $u_{0}=0$ and $f\in C^{m}([0,T];L^{2}(\Omega))$, the weak solution $u(t)$ is unique, the following prior estimates hold for positive time:
@@
The bound of $m(z)$ (see (2.9)) is an important tool in error analysis. In Lemma 2.2 of [22], Mahata and Sinha established an estimate for $|m(z)|$ under the condition $\alpha+\beta\geq 1$. This paper extends their result to a broader range $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ with any $0<\varepsilon<1$, leading to the following result.
@@
Let $m(z)$ be defined as in (2.9) with $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ for any fixed $0<\varepsilon<1$, and $\tilde{\vartheta}:=\frac{(\alpha+\beta+\varepsilon)\pi}{2(\alpha+\beta)}$. Then for any $z\in\Sigma_{\tilde{\vartheta}}$, there hold
@@
    $m(z)\in\Sigma_{\zeta}$ with $\zeta=\frac{(\alpha+\beta+\varepsilon)\pi}{2}$,
@@
Consequently, it leads to the following resolvent estimate:
$\|(m(z)I+A)^{-1}\|_{L^{2}(\Omega)}\leq C|m(z)|^{-1}\quad\forall z\in\Sigma_{\tilde{\vartheta}}$. (3.2)
@@
$|\arg(m_{1})|=\beta|\psi|<\frac{\beta}{\alpha+\beta}\cdot\frac{(\alpha+\beta+\varepsilon)\pi}{2}<\frac{(\alpha+\beta+\varepsilon)\pi}{2},$
@@
Hence, $m(z)\in\Sigma_{\zeta}$ with $\zeta=\frac{(\alpha+\beta+\varepsilon)\pi}{2}$.∎
@@
The result in Lemma 2 implies that $m(z)$ can not be negative real number. As a result, $(m(z)I+A)^{-1}$ is well-defined, and $\widehat{u}(z)$ in (2.8) only has a singular point $z=0$.
@@
Now, the determination of parameter $\tilde{c}$, a question left open in Section 2, can also be addressed. As indicated by (2.11), the slope of the asymptote of the hyperbolic contour $\Gamma$ is given by $\frac{\sin(\theta+c)}{\cos(\theta+c)}=\cot(\theta+c),$ that is, $\tan\delta=\cot(\theta+c),$ where $\delta$, as introduced earlier, denotes the angle between this asymptote and the real axis. It follows that
@@
In fact, since $\alpha+\beta\in(\varepsilon,2-\varepsilon)$, then $\frac{\varepsilon\pi}{2(\alpha+\beta)}-\theta<\frac{\pi}{2}-\theta$. Thus, based on the discussions in Subsection 2.2, it follows
$\tilde{c}=\min\Big\{\theta,\frac{\varepsilon\pi}{2(\alpha+\beta)}-\theta\Big\}.$
@@ 3.2 Error Analysis in Temporal Semi-discrete CIM Scheme
In this subsection, we analyze the error estimate for the temporal semi-discrete CIM scheme (2.12). First, it is essential to derive the decay property of $g(\phi,t)$ in (2.4) as presented in the following theorem.
@@
Let $g(\phi,t)$ be defined as in (2.4), where $z(\phi)\in S\subseteq\Sigma_{\tilde{\vartheta}}$, $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ with $0<\varepsilon<1$, and $\phi=\nu+ic\in K$, $|c|<\tilde{c}$. For $t>0$, if $u_{0}\in L^{2}(\Omega)$, and $\|\widehat{f}(z)\|_{S}<\infty$ defined in (3.1), then there exists a constant $C>0$ such that
$\|g(\phi,t)\|_{L^{2}(\Omega)}\leq C\varphi(\theta,\tilde{c})e^{\mu t}\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu},$ (3.3)
@@
Firstly, we estimate $g(\phi,t)$ on the upper part of the domain $K$: $\phi_{+}=\nu+ic\in K$, $0<c<\tilde{c}$. From (2.4) and (2.10), we obtain
$g(\nu+ic,t)=\frac{\mu}{2\pi}e^{\mu t\{1-\sin(\theta+c-i\nu)\}}\cos(\theta+c-i\nu)\widehat{u}\big(\mu\{1-\sin(\theta+c-i\nu)\}\big).$
Let $\varpi=\theta+c<\theta+\tilde{c}$. Then the following estimate holds:
$\|g(\nu+ic,t)\|_{L^{2}(\Omega)}\leq Ce^{\mu t(1-\sin\varpi\cosh\nu)}|\cos(\varpi-i\nu)|\cdot\big\|\widehat{u}\big(\mu\{1-\sin(\varpi-i\nu)\}\big)\big\|_{L^{2}(\Omega)}.$ (3.4)
The main task now is to estimate $\widehat{u}(z)$. Using the expression of $\widehat{u}(z)$ in (2.8) and the prior estimate (3.2), it follows that for all $z\in\Sigma_{\tilde{\vartheta}}$ with $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ and $0<\varepsilon<1$,
$\|\widehat{u}(z)\|_{L^{2}(\Omega)}$ $=\Big\|(m(z)I+A)^{-1}\Big(\frac{m(z)}{z}u_{0}+\frac{m(z)}{z^{\beta}}\widehat{f}(z)\Big)\Big\|_{L^{2}(\Omega)}$
@@
$|\cos(\varpi-i\nu)|\cdot\big\|\widehat{u}\big(\mu\{1-\sin(\varpi-i\nu)\}\big)\big\|_{L^{2}(\Omega)}$ (3.5)
$\leq C\frac{|\cos(\varpi-i\nu)|}{|\mu(1-\sin(\varpi-i\nu))|}\|u_{0}\|_{L^{2}(\Omega)}+C\frac{|\cos(\varpi-i\nu)|}{|\mu(1-\sin(\varpi-i\nu))|^{\beta}}\|\widehat{f}(z)\|_{L^{2}(\Omega)}$
$\leq C\sqrt{\frac{1+\sin\varpi}{1-\sin\varpi}}\|u_{0}\|_{L^{2}(\Omega)}+C(|1-\sin(\varpi-i\nu)|^{1-\beta})\sqrt{\frac{1+\sin\varpi}{1-\sin\varpi}}\|\widehat{f}(z)\|_{L^{2}(\Omega)},$
where in the second inequality of (3.5), we employ the following result from [17]:
$\Big|\frac{\cos(\varpi-i\nu)}{(1-\sin(\varpi-i\nu))}\Big|^{2}=\frac{\cosh\nu+\sin\varpi}{\cosh\nu-\sin\varpi}\leq\frac{1+\sin\varpi}{1-\sin\varpi}.$
Given that $0<\beta<1$, it follows that $|1-\sin(\varpi-i\nu)|^{1-\beta}\leq 2^{1-\beta}$. Then according to (3.4) and (3.5), there holds
$\|g(\nu+ic,t)\|_{L^{2}(\Omega)}\leq C\sqrt{\frac{1+\sin(\theta+\tilde{c})}{1-\sin(\theta+\tilde{c})}}e^{\mu t}\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu}.$
@@
$\|g(\nu-ic,t)\|_{L^{2}(\Omega)}\leq C\sqrt{\frac{1+\sin(\theta+\tilde{c})}{1-\sin(\theta+\tilde{c})}}e^{\mu t}\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu}.$
@@
From Theorem 1, it can be observed that $g(\nu+ic,t)$ exhibits biexponential decay w.r.t $\nu$, that is, it decays quite rapidly as $\phi=\nu+ic$, $\nu\to\pm\infty$. To further analyze the error of the temporal semi-discrete CIM scheme (2.12), we need to introduce the following lemma.
@@
$\int_{l}^{+\infty}e^{-\gamma\cosh x}\,dx\leq\big(1+L(\gamma)\big)e^{-\gamma\cosh l},\quad l>0,~\gamma>0.$
As noted in [39], each new value of $t$ leads to different parameters $\theta$ and $\tilde{c}$, implying that the transform $\widehat{u}(z)$ must be sampled on a distinct contour. This could result in computational inefficiency in the matrix problem associated with the discretization of the spatial operator $A$ in (2.15), as well as in the observation of the long-term dynamic evolution of the simulated solution over time. To address this issue and select a unified optimal contour $\Gamma$, a scaling parameter $\Lambda>1$ is introduced. The subsequent analysis is performed over the interval $t\in[t_{0},t_{1}]$, where $t_{1}=\Lambda t_{0}$, and a uniform error analysis of the temporal semi-discrete CIM scheme (2.12) is presented in the following theorem.
@@
Let $u(t)$ and $u^{N}(t)$ be the solutions given by (2.2) and (2.12), respectively. Consider the time interval $t\in[t_{0},t_{1}]$ with $t_{1}=\Lambda t_{0}$ for some $\Lambda\geq 1$. Let $\Gamma$ be the contour defined in (2.10) with parameters $\theta>0,\mu>0$, and $z\in\Sigma_{\tilde{\vartheta}}$ with $\alpha+\beta\in(\varepsilon,2-\varepsilon)$, $0<\varepsilon<1$. Then, the following error estimate holds uniformly:
$\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}\leq C\varphi(\theta,\tilde{c})L\big(\mu t_{0}\sin(\theta-\tilde{c})\big)\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1},$